OpenStudy (anonymous):

The slope of a line is 1/3. What is the slope of a line perpendicular to this line? -3 -1/3 1/3 3

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

ITS COrrect

OpenStudy (anonymous):

3

OpenStudy (anonymous):

thats -3

OpenStudy (anonymous):

-1 / (1/3)

OpenStudy (anonymous):

-3 haha

OpenStudy (anonymous):

slope of perpendicular is negative reciprocal.

OpenStudy (anonymous):

To find the slope of a line perpendicular to another line, you take the negative reciprocal. So, if the slope is \(x\), you do \(\dfrac{-1}{x}\) Here, \(x=1/3\), so you do \(\dfrac{-1}{1/3}\), which comes out to -3.